Find the number of spanning trees of K3,3 using Kirchoffs matrix-tree theorem.
Answer #1

K33 is de Now the Adjaceney matrix Adjaceney matrix of the graph KB3 is. a, a b c d e f blooo 11 0 0 1 1 А | op el 0 al - - -1 -2 NN 3 o = (-1)(-3) . омм wow T7 -3 -3 - 3 - 2 3 w 0 1 T 15 -3 7 -3 o -3 | Rg + R3 = 3(-3) { (-2)*3) - 5x3} = (-9) 46-15)